# Pseudo-Legendrian and Legendrian Simplicity of Links in 3-Manifolds

### v1

@@ 1. Introduction
There has been much recent work on classification problems for Legendrian links, with many exciting new questions laid out in the work of Dalton-Etnyre-Traynor [dalton2021legendrian]. The second author and Etnyre–Min–Rodewald showed the existence of stabilized Legendrian links which are topologically isotopic and component-wise Legendrian isotopic, but not Legendrian isotopic as links [chatterjee2025cable]. The first example with non-stabilized components was observed in [jordan2006generating], where the authors used generating functions to distinguish such links. Note that all of these results are in tight contact manifolds. This paper gives similar phenomena in general overtwisted manifolds for links whose components are loose. This is surprising as overtwisted manifolds and loose knots and links are known to be more flexible than tight manifolds; on the other hand, this flexibility allows us to construct examples exhibiting non-simplicity phenomena that have not yet been studied in tight manifolds. The foundational works of Eliashberg [eliashberg1989classification, eliashberg1992contact] introduced tight and overtwisted contact manifolds and first studied the flexibility phenomenon for overtwisted contact structures.
@@
The study of Legendrian-simple and non-simple isotopy classes of knots (rather than links), in terms of generalized classical invariants, was carried out by the first and third author in [CC]. In particular, the authors constructed infinite families of Legendrian-simple and non-simple isotopy classes of knots in overtwisted contact structures $S^{1}$-bundles over surfaces, and gave sufficient conditions for an isotopy class to be Legendrian simple in an arbitrary co-oriented contact manifold $(M,\xi).$ The examples in [CC] are constructed using homotopy-theoretic data from the space of loops in $M$, and the homotopy class of $\xi$; this homotopy-theoretic answer is more subtle than one might expect. The goal of this article is to establish similar machinery for links. While theorems about the conditions under which classes are simple generalize straightforwardly from the knot setting to links, the construction of various types of non-simple classes is richer in the setting of links.
To establish the above machinery, we consider a more general problem. Let $(M,V)$ denote a compact, oriented smooth 3-manifold $M$ equipped with a nowhere-vanishing smooth vector field $V$. When $M$ is equipped with a co-oriented contact structure, we will take $V$ to be a co-orienting vector field for $\xi$. Next we consider a version of the simplicity problem for links in $(M,V)$ which are everywhere transverse to $V$; that is, the tangent vector to the link and the vector $V$ at the corresponding point span a 2-plane. Such links are called $V$-transverse. When $V$ is a co-orienting vector field for a contact structure, $V$-transverse links are also called pseudo-Legendrian. Pseudo-Legendrian links were introduced by Benedetti and Petronio [benedetti2001combed, benedetti2001reidemeister]. Each $V$-transverse link is equipped with a natural framing. In the $V$-transverse category, a framed isotopy class of links is simple if any two $V$-transverse links in that isotopy class which are homotopic as $V$-transverse immersions are isotopic as $V$-transverse links. When that is not the case, the framed isotopy class is again called non-simple. When the vector field $V$ is a co-orienting vector field for a contact structure $\xi$ on $M$, we show that if a framed isotopy class in $M$ is non-simple in the $V$-transverse category, then it is non-simple in the Legendrian category.
@@
In Corollary 3, the components of the link are loose in $(M,\xi_{k})$. Again, our examples show that both of these non-simplicity phenomena can occur in a single smooth isotopy class, by controlling the number of stabilizations of each of the link components. Note that we construct an infinite family of Legendrian links which are component-wise Legendrian isotopic but not Legendrian isotopic as links. This phenomena has been observed previously in tight manifolds [jordan2006generating, chatterjee2025cable].
Finally, we state the following theorem, which gives sufficient conditions for a framed isotopy class to be simple in the $V$-transverse category. This theorem was proven for isotopy classes of knots in [CC], and generalizes immediately to links. It highlights the fact that the above examples must sit in overtwisted contact manifolds. The key ingredient in the proof of statement (3) is that the Euler class of a tight contact structure vanishes on immersed tori, as shown in [chernov2004relative, Corollary 3.10]; in the embedded case this is due to Eliashberg [eliashberg1989classification].
Theorem 4 ([CC]).
@@ 2.1. Kinks and stabilization
[CC] The $V$-transverse knots $K^{i}$ are all isotopic are framed knots.
The $V$-transverse stabilization operation is not the same as stabilization in the Legendrian category. The operations are related as follows. The stabilization of a Legendrian link component $L$ occurs in a local chart contactomorphic to $(\mathbb{R}^{3},\xi_{std})$ where $\xi_{std}=\ker(dz-ydx)$. Such a chart always exists by Darboux’s theorem. Suppose that $L_{i,j}$ denotes the Legendrian knot obtained from $L$ by adding $i$ positive and $j$ negative stabilizations. See Figure 2. As shown in [CC], $L_{i,0}$ and $(L_{0,i})^{i}$ (the superscript denoting $V$-transverse stabilization) are isotopic as $V$-transverse knots where $V=\phi_{*}^{-1}(\frac{\partial}{\partial z})$; the isotopy can be seen in the Lagrangian projection. Since $\frac{\partial}{\partial z}$ is a co-orienting vector field for $\xi_{std}$, this $V$-transverse isotopy is also a pseudo-Legendrian isotopy.
@@ 3.1. Approximating by $V$-transverse homotopies and isotopies
We recall two propositions from [CC] for knots and generalize directly to links. To simplify notation, we write the result for knots and 2-component links, but the links can have any number of components.
Proposition 9 ([CC], Lemma 3.4).
@@
Moreover, $K$ is $V$-transversely homotopic to $K^{i}$ if and only if there is a framed self-homotopy of $K$ such that the Euler class of $V^{\perp}$ evaluated on the corresponding map of $S^{1}\times S^{1}\rightarrow M$ is $2i$, as we now recall. Let $\mathcal{C}$ denote a connected component of the space of immersed curves in $M$, containing a $V$-transverse knot $K$. Next we recall the map $h_{V}:\pi_{1}(\mathcal{C},K)\rightarrow\mathbb{Z}$ from [CC] which measures the failure of the $V$-transverse knots $K$ and $K^{i}$ to be $V$-transversely homotopic. Let $\alpha]\in\pi_{1}(\mathcal{C},K)$ be a homotopy class, with $\alpha:S^{1}\rightarrow\mathcal{C}$. This map can be regarded as a map of a torus $S^{1}\times S^{1}\rightarrow M$, also denoted $\alpha$. Recall that $e_{V^{\perp}}\in H^{2}(M;\mathbb{Z})$ denotes the Euler class of the 2-plane bundle $V^{\perp}$ on $M$. Then define
@@
[CC, Lemma 4.4] Let $K$ be a $V$-transverse knot in $M$, and let $\mathcal{C}$ be the connected component of the space of immersed curves containing $K$. Let $h_{V}$ be the kink-cancelling homomorphism defined above. Then a framed self-homotopy representing a class $\alpha\in\pi_{1}(\mathcal{C},K)$ can be $C^{0}$-approximated by a $V$-transverse homotopy from $K$ to $K^{i}$ if and only if $h_{V}[\alpha]=i$. In particular, $\alpha$ can be a approximated by a $V$-transverse self-homotopy of $K$ if and only if $\alpha]\in\text{ker }h_{V}.$
@@
Figure 4. Four different types of kinks. The coordinates $(a,b)$ represent their contribution $a$ to a local rotation number and $b$ to a local writhe number of the diagram. Pairs of kinks with opposite local rotation number and opposite local writhe number can be created or cancelled by a $V$-transverse isotopy as shown in [CC].
@@
When $K$ is equipped with a framing, we can consider the connected component of the space of framed curves $\mathcal{C}_{f}$ containing $K$. Let $\mathcal{C}_{f}^{\prime}$ be a quotient of $\mathcal{C}_{f}$, obtained by identifying pairs of framed curves whose underlying immersed curves are equal and whose framings are homotopic as nonzero sections of the induced normal bundle of $K$. There is a covering map $j:\mathcal{C}_{f}^{\prime}\rightarrow\mathcal{C}$ with discrete fiber [chernov2005framed, Section 3.1].
When $K$ is framed, each of the above loops $\alpha$, with $\alpha]\in\pi_{1}(\mathcal{C},K)$, then lifts to a path $\tilde{\alpha}$ in $\mathcal{C}_{f}^{\prime}$ based at $K$. Given any choice of framing on the basepoint curve $K$, the loop $\gamma_{s}$ (which is a self-homotopy, but not a self-isotopy of $K$) induces a full twist of the framing of $K$. Hence $\gamma_{s}$ lifts to a path, but not a loop, in $\mathcal{C}_{f}^{\prime}$. The main result of [cahn2014number], in the case where the ambient manifold is orientable, can be phrased as follows: If $\alpha$ is a self-isotopy of $K$ in $\mathcal{C}$, then the lift $\tilde{\alpha}$ is a loop in $\mathcal{C}_{f}$ if and only if $K$ does not intersect a non-separating 2-sphere in $M$. When $M$ is an $S^{1}$-bundle over an orientable surface of genus $\geq 1$, no such $2$-sphere exists. Therefore, $\gamma_{\rho}$, $\gamma_{\text{rot}}$, and $\gamma_{\text{fib}}$ all lift to loops in $\mathcal{C}_{f}$.
@@
It follows that if $\alpha_{1},\alpha_{2}\in\pi_{1}(\mathcal{C},K)$ have the same trace (where the trace is considered as an element of $\pi_{1}(STM)$), then $\alpha_{2}=\alpha_{1}\gamma_{s}^{m}$; see [chernov2005framed, Lemma 3.6]. We next show that for any $\alpha\in\pi_{1}(\mathcal{C},K)$, there exists $i\in\mathbb{Z}$ such that
@@
Let $\text{pr}^{\prime}:M\rightarrow F$ be the bundle map realizing $M$ as an $S^{1}$-fibration over $F$, with $S^{1}$-fiber $f$. We now have an abelian subgroup $\text{pr}_{*}^{\prime}(A)\leq\pi_{1}(F,p)$, which, since the genus of $F$ is at least two, lies in a unique infinite cyclic subgroup [doCarmo1992, Theorem 3.2]. Choose a generator $s$ of this subgroup and choose $\sigma\in\pi_{1}(M,K(t_{0}))$ such that $\text{pr}_{*}^{\prime}(\sigma)=s$. The long exact sequence of the fibration $\text{pr}^{\prime}:M\rightarrow F$ yields that two such choices of $\sigma$ differ by a power of $[f]$, and for any such choice of $\sigma$, $A=\langle[K],[\text{tr}(\alpha)]\rangle\leq\langle[f],\sigma\rangle.$ Hence we can write $K=\sigma^{i}[f]^{j}$ and $\text{tr}(\alpha)=\sigma^{k}[f]^{l}.$
@@ 4. A Finite-Type Invariant of $V$-transverse links
We define a function on certain generic homotopies of 2-component $V$-transverse links. Under certain conditions on the homotopy classes of the link components, this function can be used to define a finite-type invariant of $V$-transverse links analogous to invariants of smooth knots and links defined in [Vassiliev], [Kirk-Livingston], and [kalfagianni1998finite].
@@ 5. Examples
Following Arnold, it remains to check that $\nu$ vanishes on small loops around the codimension 2 stratum in $\mathcal{C}_{V}$ as in [arnold1994topological]. The only codimension 2 strata that arise are formed by $V$-transverse curves with exactly two double points. When going around the stratum, we see four singular curves with one double point separated into pairs with opposite signs, so $\nu$ indeed vanishes when going around this stratum. For a similar argument, see [tchernov2002arnold, Theorem 4.3.1 and Remark 4.2.2]. ∎
@@
According to [CC, Thoeorem 10.5], there exists no $V_{k}$-transverse isotopy between $K_{1}$ and $K_{1}^{i_{1}}$ for $i_{1}\neq 0.$ The links $K_{1}\sqcup K_{2}$ and $K_{1}^{i_{1}}\sqcup K_{2}$ are not component-wise $V_{k}$-transversely isotopic, and hence are not $V_{k}$-transversely isotopic as links.
@@
We begin stating a collection of results relating homotopy and isotopy in the $V$-transverse and Legendrian categories. A version of $(4)$ is formulated for $V$-transverse knots in [CC], is formulated in the formal Legendrian category in [cieliebak2012stein], as well as proven in other cases in increasing generality in [dymara2001legendrian, eliashberg2009topologically, ding2004surgery].
@@
    If there is an overtwisted disk $D$ in $(M,\xi)$ such that $L_{1},L_{2}\subseteq M\setminus D$, then $L_{1}$ and $L_{2}$ are component-wise isotopic as $V$-transverse links if and only if $L_{1}$ and $L_{2}$ are component-wise isotopic as Legendrian links. [cieliebak2012stein, Theorem 7.19]
@@
    If there is an overtwisted disk $D$ in $(M,\xi)$ such that $L_{1},L_{2}\subseteq M\setminus D$, then $L_{1}$ and $L_{2}$ are isotopic as $V$-transverse links if and only if $L_{1}$ and $L_{2}$ are isotopic as Legendrian links. [cieliebak2012stein, Theorem 7.19]
@@
$(1)$: Let $L_{1}=A_{1}\sqcup A_{2}$ $L_{2}=B_{1}\sqcup B_{2}$ be Legendrian links in $(M,\xi)$ which are homotopic as $V$-transverse multi-curves. After sufficiently many stabilizations of the components of $L_{1}$ and $L_{2}$, any homotopy from $L_{1}$ and $L_{2}$ carrying $A_{1}$ to $B_{1}$ and $A_{2}$ to $B_{2}$ can be $C^{0}$-approximated by a Legendrian homotopy [gromov1986partial, eliashberg1992contact]. Recall from Section 2.1 and Figure  2 that $L_{i,j}$ denotes a Legendrian link component $L$ with $i$ positive and $j$ negative stabilizations. In particular, for fixed $i=1,2$ (we repeat the argument once for each component), for any $n_{1},n_{2}\in\mathbb{Z}^{+}$, there exist $n_{3},n_{4}\in\mathbb{Z}^{+}$ such that $(A_{i})_{n_{1},n_{2}}$ is Legendrian homotopic to $(B_{i})_{n_{3},n_{4}}$, and this Legendrian homotopy is $C^{0}$-close to the given $V$-transverse homotopy from $A_{i}$ to $B_{i}$; see also [fuchs1995invariants, goryunov2001plane, hill1997vassiliev, tchernov2003isomorphism]. Let $(A_{i})_{t}$ the homotopy from $A_{i}$ to $B_{i}$ at time $t$, and similarly let $((A_{i})_{n_{1},n_{2}})_{t}$ denote the homotopy from $(A_{i})_{n_{1},n_{2}}$ to $(B_{i})_{n_{3},n_{4}}$ at time $t$. By comparing the framing of $(A_{i})_{t}$ to the framing of $((A_{i})_{n_{1},n_{2}})_{t}$ throughout the homotopy, one can show that $n_{1}+n_{2}=n_{3}+n_{4}$; by comparing the homotopy classes of the lifts of $(A_{i})_{t}$ and $((A_{i})_{n_{1},n_{2}})_{t}$ to $E_{V}$, one can show that $n_{1}-n_{2}=n_{3}-n_{4}$. Details of this argument are given in [tchernov2003isomorphism, Section 5.3]. It follows that $n_{1}=n_{3}$ and $n_{2}=n_{4}.$ Taking $n=\text{max}(n_{1},n_{2})$, we can perform extra stabilizations on the initial and final components to get a Legendrian homotopy from $(A_{i})_{n,n}$ to $(B_{i})_{n,n}$. As shown in [tchernov2003isomorphism, Figure 6], any Legendrian curve $L$ is Legendrian homotopic to $L_{n,n}$ for all $n$; this homotopy is performed locally on an arc of $L$. Hence for each fixed $i=1,2$ we get a Legendrian homotopy from $A_{i}$ to $B_{i}$, so $L_{1}$ and $L_{2}$ are Legendrian homotopic. Conversely, if $L_{1}$ and $L_{2}$ are homotopic as Legendrian multi-curves, they are clearly $V$-transverse homotopic, as $V$ co-orients $\xi.$
@@
In the formal Legendrian category, $(3)$ and $(4)$ are proven in [cieliebak2012stein, Theorem 7.19]. In [CC], an analogous statement is proven for $V$-transverse knots by applying [cieliebak2012stein, Theorem 7.19], but the argument for $V$-transverse links is analogous. ∎
We now return to the specific case where $M$ is an oriented $S^{1}$-bundle over an oriented surface $F$ of genus $\geq 2$. Consider one of the links $K_{1}\sqcup K_{2}$ in Examples  12, or  3. As in those examples, let $d$ be a curve in $M$ such that its projection $\pi(d)$ is an essential simple closed curve on $F$ disjoint from the projections of $K_{1}$ and $K_{2}$ to $F$. By work of Lutz and Martinet, we first choose a contact structure $\xi_{k}$ on $M$ such that the Euler class $e_{\xi_{k}}$ is dual to $2k\cdot[d]\in H_{1}(M,\mathbb{Z})$, where $d$ is a curve in $M$, and $\pi(d)$ is as shown in Figures  7,  8, and  9 [lutz1971quelques, martinet2006formes]; see also [ginzburg1988closed, geiges2008introduction]. Moreover, given any Legendrian link $L_{1}\sqcup L_{2}$ in $M$, we can modify $\xi_{k}$ by performing a full Lutz along a transverse unknot in the complement of $L_{1}\sqcup L_{2}$, in such a way that the new contact structure $\xi_{k}^{\prime}$ is homotopic to $\xi_{k}$ as a 2-plane field, and such that $L_{1}\sqcup L_{2}$ is a loose link in $(M,\xi_{k}^{\prime})$, i.e., there is an overtwisted disk in its complement [geiges2008introduction, Proposition 4.5.4].

### v2

@@ 1. Introduction
There has been much recent work on classification problems for Legendrian links, with many exciting new questions laid out in the work of Dalton-Etnyre-Traynor [DET24]. The second author and Etnyre–Min–Rodewald showed the existence of stabilized Legendrian links which are topologically isotopic and component-wise Legendrian isotopic, but not Legendrian isotopic as links [CEM+25]. The first example with non-stabilized components was observed in [JT06], where the authors used generating functions to distinguish such links. Note that all of these results are in tight contact manifolds. This paper gives similar phenomena in general overtwisted manifolds for links whose components are loose. This is surprising as overtwisted manifolds and loose knots and links are known to be more flexible than tight manifolds; on the other hand, this flexibility allows us to construct examples exhibiting non-simplicity phenomena that have not yet been studied in tight manifolds. The foundational works of Eliashberg [ELI89, ELI92] introduced tight and overtwisted contact manifolds and first studied the flexibility phenomenon for overtwisted contact structures.
@@
The study of Legendrian-simple and non-simple isotopy classes of knots (rather than links), in terms of generalized classical invariants, was carried out by the first and third author in [CC20]. In particular, the authors constructed infinite families of Legendrian-simple and non-simple isotopy classes of knots in overtwisted contact structures $S^{1}$-bundles over surfaces, and gave sufficient conditions for an isotopy class to be Legendrian simple in an arbitrary co-oriented contact manifold $(M,\xi).$ The examples in [CC20] are constructed using homotopy-theoretic data from the space of loops in $M$, and the homotopy class of $\xi$; this homotopy-theoretic answer is more subtle than one might expect. The goal of this article is to establish similar machinery for links. While theorems about the conditions under which classes are simple generalize straightforwardly from the knot setting to links, the construction of various types of non-simple classes is richer in the setting of links.
To establish the above machinery, we consider a more general problem. Let $(M,V)$ denote a compact, oriented smooth 3-manifold $M$ equipped with a nowhere-vanishing smooth vector field $V$. When $M$ is equipped with a co-oriented contact structure, we will take $V$ to be a co-orienting vector field for $\xi$. Next we consider a version of the simplicity problem for links in $(M,V)$ which are everywhere transverse to $V$; that is, the tangent vector to the link and the vector $V$ at the corresponding point span a 2-plane. Such links are called $V$-transverse. When $V$ is a co-orienting vector field for a contact structure, $V$-transverse links are also called pseudo-Legendrian. Pseudo-Legendrian links were introduced by Benedetti and Petronio [BP01a, BP01b]. Each $V$-transverse link is equipped with a natural framing. In the $V$-transverse category, a framed isotopy class of links is simple if any two $V$-transverse links in that isotopy class which are homotopic as $V$-transverse immersions are isotopic as $V$-transverse links. When that is not the case, the framed isotopy class is again called non-simple. When the vector field $V$ is a co-orienting vector field for a contact structure $\xi$ on $M$, we show that if a framed isotopy class in $M$ is non-simple in the $V$-transverse category, then it is non-simple in the Legendrian category.
@@
In Corollary 3, the components of the link are loose in $(M,\xi_{k})$. Again, our examples show that both of these non-simplicity phenomena can occur in a single smooth isotopy class, by controlling the number of stabilizations of each of the link components. Note that we construct an infinite family of Legendrian links which are component-wise Legendrian isotopic but not Legendrian isotopic as links. This phenomena has been observed previously in tight manifolds [JT06, CEM+25].
Finally, we state the following theorem, which gives sufficient conditions for a framed isotopy class to be simple in the $V$-transverse category. This theorem was proven for isotopy classes of knots in [CC20], and generalizes immediately to links. It highlights the fact that the above examples must sit in overtwisted contact manifolds. The key ingredient in the proof of statement (3) is that the Euler class of a tight contact structure vanishes on smooth mappings of tori, as shown in [CHE04, Corollary 3.10]; in the embedded case this is due to Eliashberg [ELI89].
Theorem 4 ([CC20]).
@@ 2.1. Kinks and stabilization
[CC20] The $V$-transverse knots $K^{i}$ are all isotopic are framed knots.
The $V$-transverse stabilization operation is not the same as stabilization in the Legendrian category. The operations are related as follows. The stabilization of a Legendrian link component $L$ occurs in a local chart contactomorphic to $(\mathbb{R}^{3},\xi_{std})$ where $\xi_{std}=\ker(dz-ydx)$. Such a chart always exists by Darboux’s theorem. Suppose that $L_{i,j}$ denotes the Legendrian knot obtained from $L$ by adding $i$ positive and $j$ negative stabilizations. See Figure 2. As shown in [CC20], $L_{i,0}$ and $(L_{0,i})^{i}$ (the superscript denoting $V$-transverse stabilization) are isotopic as $V$-transverse knots where $V=\phi_{*}^{-1}(\frac{\partial}{\partial z})$; the isotopy can be seen in the Lagrangian projection. Since $\frac{\partial}{\partial z}$ is a co-orienting vector field for $\xi_{std}$, this $V$-transverse isotopy is also a pseudo-Legendrian isotopy.
@@ 3.1. Approximating by $V$-transverse homotopies and isotopies
We recall two propositions from [CC20] for knots and generalize directly to links. To simplify notation, we write the result for knots and 2-component links, but the links can have any number of components.
Proposition 9 ([CC20], Lemma 3.4).
@@
Moreover, $K$ is $V$-transversely homotopic to $K^{i}$ if and only if there is a framed self-homotopy of $K$ such that the Euler class of $V^{\perp}$ evaluated on the corresponding map of $S^{1}\times S^{1}\rightarrow M$ is $2i$, as we now recall. Let $\mathcal{C}$ denote a connected component of the space of immersed curves in $M$, containing a $V$-transverse knot $K$. Next we recall the map $h_{V}:\pi_{1}(\mathcal{C},K)\rightarrow\mathbb{Z}$ from [CC20] which measures the failure of the $V$-transverse knots $K$ and $K^{i}$ to be $V$-transversely homotopic. Let $\alpha]\in\pi_{1}(\mathcal{C},K)$ be a homotopy class, with $\alpha:S^{1}\rightarrow\mathcal{C}$. This map can be regarded as a map of a torus $S^{1}\times S^{1}\rightarrow M$, also denoted $\alpha$. Recall that $e_{V^{\perp}}\in H^{2}(M;\mathbb{Z})$ denotes the Euler class of the 2-plane bundle $V^{\perp}$ on $M$. Then define
@@
[CC20, Lemma 4.4] Let $K$ be a $V$-transverse knot in $M$, and let $\mathcal{C}$ be the connected component of the space of immersed curves containing $K$. Let $h_{V}$ be the kink-cancelling homomorphism defined above. Then a framed self-homotopy representing a class $\alpha\in\pi_{1}(\mathcal{C},K)$ can be $C^{0}$-approximated by a $V$-transverse homotopy from $K$ to $K^{i}$ if and only if $h_{V}[\alpha]=i$. In particular, $\alpha$ can be a approximated by a $V$-transverse self-homotopy of $K$ if and only if $\alpha]\in\text{ker }h_{V}.$
@@
Figure 4. Four different types of kinks. The coordinates $(a,b)$ represent their contribution $a$ to a local rotation number and $b$ to a local writhe number of the diagram. Pairs of kinks with opposite local rotation number and opposite local writhe number can be created or cancelled by a $V$-transverse isotopy as shown in [CC20].
@@
When $K$ is equipped with a framing, we can consider the connected component of the space of framed curves $\mathcal{C}_{f}$ containing $K$. Let $\mathcal{C}_{f}^{\prime}$ be a quotient of $\mathcal{C}_{f}$, obtained by identifying pairs of framed curves whose underlying immersed curves are equal and whose framings are homotopic as nonzero sections of the induced normal bundle of $K$. There is a covering map $j:\mathcal{C}_{f}^{\prime}\rightarrow\mathcal{C}$ with discrete fiber [CHE05, Section 3.1].
When $K$ is framed, each of the above loops $\alpha$, with $\alpha]\in\pi_{1}(\mathcal{C},K)$, then lifts to a path $\tilde{\alpha}$ in $\mathcal{C}_{f}^{\prime}$ based at $K$. Given any choice of framing on the basepoint curve $K$, the loop $\gamma_{s}$ (which is a self-homotopy, but not a self-isotopy of $K$) induces a full twist of the framing of $K$. Hence $\gamma_{s}$ lifts to a path, but not a loop, in $\mathcal{C}_{f}^{\prime}$. The main result of [CCS14], in the case where the ambient manifold is orientable, can be phrased as follows: If $\alpha$ is a self-isotopy of $K$ in $\mathcal{C}$, then the lift $\tilde{\alpha}$ is a loop in $\mathcal{C}_{f}$ if and only if $K$ does not intersect a non-separating 2-sphere in $M$. When $M$ is an $S^{1}$-bundle over an orientable surface of genus $\geq 1$, no such $2$-sphere exists. Therefore, $\gamma_{\rho}$, $\gamma_{\text{rot}}$, and $\gamma_{\text{fib}}$ all lift to loops in $\mathcal{C}_{f}$.
@@
It follows that if $\alpha_{1},\alpha_{2}\in\pi_{1}(\mathcal{C},K)$ have the same trace (where the trace is considered as an element of $\pi_{1}(STM)$), then $\alpha_{2}=\alpha_{1}\gamma_{s}^{m}$; see [CHE05, Lemma 3.6]. We next show that for any $\alpha\in\pi_{1}(\mathcal{C},K)$, there exists $i\in\mathbb{Z}$ such that
@@
Let $\text{pr}^{\prime}:M\rightarrow F$ be the bundle map realizing $M$ as an $S^{1}$-fibration over $F$, with $S^{1}$-fiber $f$. We now have an abelian subgroup $\text{pr}_{*}^{\prime}(A)\leq\pi_{1}(F,p)$, which, since the genus of $F$ is at least two, lies in a unique infinite cyclic subgroup [DO 92, Theorem 3.2]. Choose a generator $s$ of this subgroup and choose $\sigma\in\pi_{1}(M,K(t_{0}))$ such that $\text{pr}_{*}^{\prime}(\sigma)=s$. The long exact sequence of the fibration $\text{pr}^{\prime}:M\rightarrow F$ yields that two such choices of $\sigma$ differ by a power of $[f]$, and for any such choice of $\sigma$, $A=\langle[K],[\text{tr}(\alpha)]\rangle\leq\langle[f],\sigma\rangle.$ Hence we can write $K=\sigma^{i}[f]^{j}$ and $\text{tr}(\alpha)=\sigma^{k}[f]^{l}.$
@@ 4. A Finite-Type Invariant of $V$-transverse links
We define a function on certain generic homotopies of 2-component $V$-transverse links. Under certain conditions on the homotopy classes of the link components, this function can be used to define a finite-type invariant of $V$-transverse links analogous to invariants of smooth knots and links defined in [VAS90], [KL98], and [KAL98].
@@ 5. Examples
Following Arnold, it remains to check that $\nu$ vanishes on small loops around the codimension 2 stratum in $\mathcal{C}_{V}$ as in [ARN94]. The only codimension 2 strata that arise are formed by $V$-transverse curves with exactly two double points. When going around the stratum, we see four singular curves with one double point separated into pairs with opposite signs, so $\nu$ indeed vanishes when going around this stratum. For a similar argument, see [TCH02, Theorem 4.3.1 and Remark 4.2.2]. ∎
@@
According to [CC20, Thoeorem 10.5], there exists no $V_{k}$-transverse isotopy between $K_{1}$ and $K_{1}^{i_{1}}$ for $i_{1}\neq 0.$ The links $K_{1}\sqcup K_{2}$ and $K_{1}^{i_{1}}\sqcup K_{2}$ are not component-wise $V_{k}$-transversely isotopic, and hence are not $V_{k}$-transversely isotopic as links.
@@
We begin stating a collection of results relating homotopy and isotopy in the $V$-transverse and Legendrian categories. A version of $(4)$ is formulated for $V$-transverse knots in [CC20], is formulated in the formal Legendrian category in [CE12], as well as proven in other cases in increasing generality in [DYM01, EF09, DGS04].
@@
    If there is an overtwisted disk $D$ in $(M,\xi)$ such that $L_{1},L_{2}\subseteq M\setminus D$, then $L_{1}$ and $L_{2}$ are component-wise isotopic as $V$-transverse links if and only if $L_{1}$ and $L_{2}$ are component-wise isotopic as Legendrian links. [CE12, Theorem 7.19]
@@
    If there is an overtwisted disk $D$ in $(M,\xi)$ such that $L_{1},L_{2}\subseteq M\setminus D$, then $L_{1}$ and $L_{2}$ are isotopic as $V$-transverse links if and only if $L_{1}$ and $L_{2}$ are isotopic as Legendrian links. [CE12, Theorem 7.19]
@@
$(1)$: Let $L_{1}=A_{1}\sqcup A_{2}$ $L_{2}=B_{1}\sqcup B_{2}$ be Legendrian links in $(M,\xi)$ which are homotopic as $V$-transverse multi-curves. After sufficiently many stabilizations of the components of $L_{1}$ and $L_{2}$, any homotopy from $L_{1}$ and $L_{2}$ carrying $A_{1}$ to $B_{1}$ and $A_{2}$ to $B_{2}$ can be $C^{0}$-approximated by a Legendrian homotopy [GRO86, ELI92]. Recall from Section 2.1 and Figure  2 that $L_{i,j}$ denotes a Legendrian link component $L$ with $i$ positive and $j$ negative stabilizations. In particular, for fixed $i=1,2$ (we repeat the argument once for each component), for any $n_{1},n_{2}\in\mathbb{Z}^{+}$, there exist $n_{3},n_{4}\in\mathbb{Z}^{+}$ such that $(A_{i})_{n_{1},n_{2}}$ is Legendrian homotopic to $(B_{i})_{n_{3},n_{4}}$, and this Legendrian homotopy is $C^{0}$-close to the given $V$-transverse homotopy from $A_{i}$ to $B_{i}$; see also [FT95, GOR01, HIL97, TCH03]. Let $(A_{i})_{t}$ the homotopy from $A_{i}$ to $B_{i}$ at time $t$, and similarly let $((A_{i})_{n_{1},n_{2}})_{t}$ denote the homotopy from $(A_{i})_{n_{1},n_{2}}$ to $(B_{i})_{n_{3},n_{4}}$ at time $t$. By comparing the framing of $(A_{i})_{t}$ to the framing of $((A_{i})_{n_{1},n_{2}})_{t}$ throughout the homotopy, one can show that $n_{1}+n_{2}=n_{3}+n_{4}$; by comparing the homotopy classes of the lifts of $(A_{i})_{t}$ and $((A_{i})_{n_{1},n_{2}})_{t}$ to $E_{V}$, one can show that $n_{1}-n_{2}=n_{3}-n_{4}$. Details of this argument are given in [TCH03, Section 5.3]. It follows that $n_{1}=n_{3}$ and $n_{2}=n_{4}.$ Taking $n=\text{max}(n_{1},n_{2})$, we can perform extra stabilizations on the initial and final components to get a Legendrian homotopy from $(A_{i})_{n,n}$ to $(B_{i})_{n,n}$. As shown in [TCH03, Figure 6], any Legendrian curve $L$ is Legendrian homotopic to $L_{n,n}$ for all $n$; this homotopy is performed locally on an arc of $L$. Hence for each fixed $i=1,2$ we get a Legendrian homotopy from $A_{i}$ to $B_{i}$, so $L_{1}$ and $L_{2}$ are Legendrian homotopic. Conversely, if $L_{1}$ and $L_{2}$ are homotopic as Legendrian multi-curves, they are clearly $V$-transverse homotopic, as $V$ co-orients $\xi.$
@@
In the formal Legendrian category, $(3)$ and $(4)$ are proven in [CE12, Theorem 7.19]. In [CC20], an analogous statement is proven for $V$-transverse knots by applying [CE12, Theorem 7.19], but the argument for $V$-transverse links is analogous. ∎
We now return to the specific case where $M$ is an oriented $S^{1}$-bundle over an oriented surface $F$ of genus $\geq 2$. Consider one of the links $K_{1}\sqcup K_{2}$ in Examples  12, or  3. As in those examples, let $d$ be a curve in $M$ such that its projection $\pi(d)$ is an essential simple closed curve on $F$ disjoint from the projections of $K_{1}$ and $K_{2}$ to $F$. By work of Lutz and Martinet, we first choose a contact structure $\xi_{k}$ on $M$ such that the Euler class $e_{\xi_{k}}$ is dual to $2k\cdot[d]\in H_{1}(M,\mathbb{Z})$, where $d$ is a curve in $M$, and $\pi(d)$ is as shown in Figures  7,  8, and  9 [LUT71, MAR06]; see also [GIN88, GEI08]. Moreover, given any Legendrian link $L_{1}\sqcup L_{2}$ in $M$, we can modify $\xi_{k}$ by performing a full Lutz along a transverse unknot in the complement of $L_{1}\sqcup L_{2}$, in such a way that the new contact structure $\xi_{k}^{\prime}$ is homotopic to $\xi_{k}$ as a 2-plane field, and such that $L_{1}\sqcup L_{2}$ is a loose link in $(M,\xi_{k}^{\prime})$, i.e., there is an overtwisted disk in its complement [GEI08, Proposition 4.5.4].